# Transfer Learning for High-dimensional Quantile Regression with Distribution Shift

### v1

@@ 4.1 Debiased Estimator based on Neyman Orthogonality
where $\lambda_{\boldsymbol{\theta}}^{(k)}$ is a tuning parameter and its value will be shown in the simulation section. For each high-dimensional nuisance parameter $\boldsymbol{w}_{-1}^{(k)}$, we calculate the post-selected QR estimator (denoted by $\widetilde{\boldsymbol{w}}_{-1}^{(k)}$), since the refitting step after feature screening can lead to better finite sample behaviors (Belloni et al.,, 2019). Specifically, we screen each source study’s covariates by setting threshold for each element of $\widehat{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)}$ obtained in Algorithm 2, then apply traditional QR on the selected features. The empirical score function corresponding to (11) is consequently defined as
@@
$\widehat{\psi}_{i}^{(k)}(\alpha,\widetilde{\boldsymbol{w}}_{-1}^{(k)},\widehat{\boldsymbol{\theta}}^{(k)},\widehat{f}_{i}^{(k)}):=\left(\mathbb{I}\left\{y_{i}^{(k)}\leq x_{i,1}^{(k)}\alpha+(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\widetilde{\boldsymbol{w}}^{(k)}_{-1}\right\}-\tau\right)\left[\widehat{f}^{(k)}_{i}x_{i,1}^{(k)}-\widehat{f}^{(k)}_{i}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\widehat{\boldsymbol{\theta}}^{(k)}\right].$ (15)
Finally, following the moment condition (12), we aggregate the empirical score functions within the set $\widehat{\mathcal{C}}$ from Algorithm 1 and compute the debiased estimate $\widehat{\alpha}$ for $\beta_{1}^{*}$ by
$\underset{\alpha\in\mathcal{A}_{\widehat{\mathcal{C}}}}{\arg\min}\left|\sum_{k\in\{0\}\cup\widehat{\mathcal{C}}}\sum_{i=1}^{n_{k}}\widehat{\psi}_{i}^{(k)}(\alpha,\widetilde{\boldsymbol{w}}_{-1}^{(k)},\widehat{\boldsymbol{\theta}}^{(k)},\widehat{f}_{i}^{(k)})\right|,$
@@
where $\mathcal{A}_{\widehat{\mathcal{C}}}=\{\alpha\in\mathbb{R}:|\alpha-\widehat{\beta}_{1}|\leq 10[n_{\widehat{\mathcal{C}}}^{-1}\sum_{k\in\{0\}\cup\widehat{\mathcal{C}}}\sum_{i=1}^{n_{k}}(x^{(k)}_{i,1})^{2}]^{-1/2}/\log n_{\widehat{\mathcal{C}}}\}$ is the search region for theoretical consideration, as suggested by Belloni et al., (2019). The whole inference procedure is summarized in Algorithm 3.
Tackling the trouble by approximation error $r_{i}^{(k)}$. As noted by Belloni et al., (2019), the orthogonal approach remains theoretically valid as long as $\mathbb{E}[(r_{i}^{(k)})^{2}]=(\mathbb{E}[x_{i1}^{(k)}]\delta_{1}^{(k)})^{2}$ stays below a certain threshold. In the transfer setting, introducing more external data increases the requirement for the contrast $\delta_{1}^{(k)}$, where a sufficient theoretical bound is $o(n_{\widehat{\mathcal{C}}}^{-1/2})$. However, constructing consistent selection procedures or hypothesis tests for this condition is nearly impossible, especially when dynamically selecting $\widehat{\mathcal{C}}$. Although there is no theoretical guarantee, we propose a practical quality control procedure to validate the transfer debiased estimator. We can use prior knowledge to exclude certain source studies with non-negligible $\widetilde{\delta}_{1}^{(k)}$ and perform a bootstrap test to check the normality of the debiased estimator. If the normality hypothesis is rejected, we revert to debiasing only the target study, yielding a result similar to previous approaches. The full procedure is outlined in Algorithm 4 in the supplementary materials.
@@
Input: Detected transferable set $\widehat{\mathcal{C}}$ and estimators $\widehat{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k\in\widehat{\mathcal{C}}}$ from Algorithm 2, target study $\{((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}\}$, source studies $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k\in\widehat{\mathcal{C}}}$, tuning parameters $\{\lambda^{(k)},\lambda_{\boldsymbol{\theta}}^{(k)}\}_{k\in\{0\}\cup\widehat{\mathcal{C}}}$, threshold parameter $\bar{\lambda}$.
@@
    For each study $k\in\{0\}\cup\widehat{\mathcal{C}}$, estimate the $\tau$-th densities $\widehat{f}_{i}^{(k)}$ by (13), conduct the projection (14), and calculate the residuals $\widehat{v}_{i}^{(k)}=\widehat{f}^{(k)}_{i}x_{i,1}^{(k)}-\widehat{f}^{(k)}_{i}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\widehat{\boldsymbol{\theta}}^{(k)}$.
@@
    For each study $k\in\{0\}\cup\widehat{\mathcal{C}}$, compute $\widehat{\boldsymbol{w}}^{(k)}$ by traditional QR on selected data $\{((\boldsymbol{x}_{i,\widehat{S}^{(k)}}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}$, where the index set $\widehat{S}^{(k)}=\{j:|\widehat{\beta}_{j}+\widetilde{\delta}^{(k)}_{j}|>\bar{\lambda}\}$, then enlarge $\widehat{\boldsymbol{w}}^{(k)}$ (with zero) to the $p$-dimensional vector $\widetilde{\boldsymbol{w}}^{(k)}$.
@@
    Incorporate empirical score functions (15) within target and transferable source studies and solve for
@@ 4.2 Asymptotic Normality for Debiased Estimator
Before examining the asymptotic normality of the estimator $\widehat{\alpha}$ in Algorithm 3, we introduce regularity conditions that ensure the validity of the inference results. As in Theorem 1, we assume an arbitrarily specified transferable set $\mathcal{C}$. Let $c$ and $C$ denote given positive constants. Although these could vary across studies, incorporating them into the growth conditions in Theorem 3 below would complicate the analysis significantly. Therefore, for simplicity, we treat them as universal fixed values.
@@
For each study $k\in\{0\}\cup\mathcal{C}$, (i) assume each $\|\boldsymbol{\theta}_{0}^{(k)}\|_{2}\leq C$ (ii) There exists a universal sparsity level $s^{*}$ and vectors $\{\boldsymbol{\theta}^{(k)}\}_{k\in\{0\}\cup\mathcal{C}}$ such that $s^{*}=\max_{k\in\{0\}\cup\mathcal{C}}\{s+s_{k}+\|\boldsymbol{\theta}^{(k)}\|_{0}\}$, where each $\boldsymbol{\theta}^{(k)}$ satisfies $(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{\theta}_{0}^{(k)}=(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{\theta}^{(k)}+r_{i\boldsymbol{\theta}}^{(k)}$ with $\|\boldsymbol{\theta}_{0}^{(k)}-\boldsymbol{\theta}^{(k)}\|_{1}=Cs^{*}\sqrt{\log p/n_{k}}$, $\mathbb{E}[(r_{i\boldsymbol{\theta}}^{(k)})^{2}]\leq Cs^{*}/n_{k}$. (iii) We have $c\leq\min_{1<j\leq p}\mathbb{E}[(f_{i}^{(k)}x^{(k)}_{i,j}v_{0i}^{(k)})^{2}]^{1/2}\leq\max_{1<j\leq p}\mathbb{E}[(f_{i}^{(k)}x^{(k)}_{i,j}v_{0i}^{(k)})^{3}]^{1/3}\leq C$.
@@
For each study $k\in\{0\}\cup\mathcal{C}$, (i) conditional density $f(\cdot|\boldsymbol{x}_{i}^{(k)})$ are continuously differentiable and uniformly lower bounded by a positive constant $\underline{f}$. (ii) For $u=\tau\pm h_{k}$, assume that $u$-quantile$(y_{i}^{(k)}|\boldsymbol{x}_{i}^{(k)})=(\boldsymbol{x}_{i}^{(k)})^{\top}\boldsymbol{w}^{(k)}_{u}+\epsilon_{iu}^{(k)}$, where $\mathbb{E}[\epsilon_{iu}^{(k)}]\leq o(n_{k}^{-1/2}),|\epsilon_{iu}^{(k)}|\leq o(h_{k})$ and $|\boldsymbol{w}^{(k)}_{u}|_{0}\leq s^{*}$.
Condition 7-9 actually generalize the assumptions in Belloni et al., (2019) to the transfer setting. Condition 7 follows the original settings, providing the necessary requirements for establishing convergence rates of weighted Lasso estimators $\{\widehat{\boldsymbol{\theta}}^{(k)}\}_{k\in\{0\}\cup\mathcal{C}}$. Condition 7 (ii) is to assume the projection parameter $\boldsymbol{\theta}^{(k)}_{0}$ in (10) can be approximated by the sparse vector $\boldsymbol{\theta}^{(k)}$ up to a small error. while Condition 7 (iii) imposes necessary mild moment conditions on the variables $f_{i}^{(k)}\boldsymbol{x}^{(k)}_{i,j}v_{0i}^{(k)}$. Condition 8 set up the moment requirements for approximation error $r_{i}^{(k)}$. Since the debiased sample size $n_{\mathcal{C}}$ is larger than the target size $n_{0}$, Condition 8 (ii) and (iii) have to link each source’s quantity $r_{i}^{k}$ to the whole sample size $n_{\mathcal{C}}$. Admittedly, due to the need of approximate orthogonality, each $\mathbb{E}[f_{i}^{(k)}v_{0i}^{(k)}r_{i}^{(k)}]=\mathbb{E}[f_{i}^{(k)}x_{i1}^{(k)}v_{0i}^{(k)}\delta_{1}^{(k)}]=\mathbb{E}[(f_{i}^{(k)}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\boldsymbol{\theta}_{0}^{(k)}+v_{0i}^{(k)})v_{0i}^{(k)}\delta_{1}^{(k)}]=\mathbb{E}[(v_{0i}^{(k)})^{2}]\delta_{1}^{(k)}$ needs to very small, where a sufficient condition is $\delta^{(k)}_{1}=o(n_{\mathcal{C}}^{-1/2})$. This requirement is understandable since we need smaller parameter contrast for better normality results, plus we don’t have any approaches to remove this error. Condition 9 is for the estimation of the conditional density at $(\tau\pm h_{k})$-th quantiles, where the approximately sparse QR model is assumed around the quantile index $\tau$. All of these conditions follow the corresponding version of single-study case established in Belloni et al., (2019).
@@
Under Condition 1-9, if the parameters satisfy the following growth conditions
$\frac{s\log p}{n_{\mathcal{C}}}+\sum_{k\in\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}\|\boldsymbol{\delta}^{(k)}\|_{1}=o(1/\sqrt{n_{\mathcal{C}}}),$
$\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\left[\frac{1}{h_{k}}\sqrt{\frac{s^{*}\log p}{n_{k}}}+h_{k}^{2}+\frac{\lambda_{\boldsymbol{\theta}}^{(k)}s^{*}}{n_{k}}\right]=o(n_{\mathcal{C}}^{-1/4}),$
$\max_{k\in\{0\}\cup\mathcal{C}}\left\{s^{*}+\frac{n_{k}s^{*}\log p}{(h_{k}\lambda^{(k)}_{\boldsymbol{\theta}})^{2}}+\left(\frac{n_{k}h_{k}^{2}}{\lambda^{(k)}_{\boldsymbol{\theta}}}\right)^{2}\right\}\frac{\log p}{n_{\mathcal{C}}}=o(1/\sqrt{n_{\mathcal{C}}}).$
Then our estimator $\widehat{\alpha}$ based on Algorithm 3 satisfies
$\qquad\qquad\qquad\sigma_{\mathcal{C}}^{-1}\sqrt{n_{\mathcal{C}}}(\widehat{\alpha}-\beta_{1}^{*})=\mathbb{U}_{\mathcal{C}}(\tau)+o_{p}(1),\qquad\mathbb{U}_{\mathcal{C}}(\tau)\leadsto\mathcal{N}(0,1),$
$\text{where}\quad\mathbb{U}_{\mathcal{C}}(\tau)=\left(\tau(1-\tau)\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[(v_{0i}^{(k)})^{2}]\right)^{-1/2}\cdot\frac{1}{\sqrt{n_{\mathcal{C}}}}\sum_{k\in\{0\}\cup\mathcal{C}}\sum_{i=1}^{n_{k}}\psi_{i}^{(k)}(\alpha,\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)}),$
$\text{and}\quad\sigma_{\mathcal{C}}^{2}=\left(\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[f_{i}^{(k)}x_{i,1}^{(k)}v_{0i}^{(k)}]\right)^{-2}\cdot\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[\tau(1-\tau)(v_{0i}^{(k)})^{2}].$
@@
With the asymptotic normality of $\widehat{\alpha}$ and consistent variance estimators $\widehat{\sigma}_{\mathcal{C}}^{2}$ provided in Theorem 3, we can construct valid confidence intervals and hypothesis test procedures for the objective component $\beta^{*}_{1}$. Now we give some explanations for results in Theorem 3. Firstly, our estimator $\widehat{\alpha}$ enjoys $\sqrt{n_{\mathcal{C}}}$-normality, which is much sharper than the previous results, demonstrating the effectiveness of our debiased transfer methods. Secondly, conditions of Theorem 3 allow for nonidentical covariate distributions across studies, as well as robust to the divergence of covariance matrices between target and source studies, demonstrating our methods’ superiority under covariate shift. Moreover, note that the asymptotic variance $\sigma_{\mathcal{C}}^{2}$ includes $v_{0i}^{(k)}$ that is inversely proportional to the $\tau$-th density $f_{i}^{(k)}$, clearly revealing that residual shift may increase the variance and deteriorate the inference efficiency, and that’s why we choose to only debias on the detected transferable set in Algorithm 3. Although the transferable set and orthogonal technique can alleviate the bias a lot, leveraging source data still needs parameter shift small enough, especially for the approximation error $x_{i1}^{(k)}\delta_{1}^{(k)}$. In the following simulations we will show that transfer debiased estimators can enjoy sharper inference results if the contrast level $\delta_{1}^{(k)}$ is small, while may lose power when higher parameter shift is introduced.

### v2

@@ 4.1 Debiased Estimator based on Neyman Orthogonality
For each high-dimensional nuisance parameter $\boldsymbol{w}_{-1}^{(k)}$, we calculate the post-selected QR estimator (denoted by $\widetilde{\boldsymbol{w}}_{-1}^{(k)}$), since the refitting step after feature screening can lead to better finite sample behaviors [belloni2019valid]. Specifically, we screen each source study’s covariates by setting threshold for each element of $\widehat{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)}$ obtained in Algorithm 1, then apply traditional QR on the selected features. Denote the empirical residuals $\widehat{v}_{i}^{(k)}=\widehat{f}^{(k)}_{i}x_{i,1}^{(k)}-\widehat{f}^{(k)}_{i}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\widehat{\boldsymbol{\theta}}^{(k)}$. The empirical score function corresponding to (10) is defined as
@@
$\widehat{\psi}_{i}^{(k)}(\alpha,\widetilde{\boldsymbol{w}}_{-1}^{(k)},\widehat{\boldsymbol{\theta}}^{(k)},\widehat{f}_{i}^{(k)}):=\left[\mathbb{I}\left\{y_{i}^{(k)}\leq x_{i,1}^{(k)}\alpha+(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\widetilde{\boldsymbol{w}}^{(k)}_{-1}\right\}-\tau\right]\widehat{v}_{i}^{(k)}.$ (12)
@@
To tackle parameter and residual shift, we only incorporate the score functions of sources within the set $\widehat{\mathcal{C}}$ from Algorithm 1. Finally, we set the search region $\widehat{\alpha}\in\mathcal{A}_{\widehat{\mathcal{C}}}:=\{\alpha\in\mathbb{R}:|\alpha-\widehat{\beta}_{1}|\leq 10[n_{\widehat{\mathcal{C}}}^{-1}\sum_{k\in\{0\}\cup\widehat{\mathcal{C}}}\sum_{i=1}^{n_{k}}(x^{(k)}_{i,1})^{2}]^{-1/2}/\log n_{\widehat{\mathcal{C}}}\}$, as suggested by [belloni2019valid]. The whole inference procedure is summarized in Algorithm 2.
@@
Input: Detected transferable set $\widehat{\mathcal{C}}$ and estimators $\widehat{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k\in\widehat{\mathcal{C}}}$ from Algorithm 1,    target study $\{((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}\}$, source studies $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k\in\widehat{\mathcal{C}}}$,    tuning parameters $\{\lambda_{\tau}^{(k)},\lambda_{\boldsymbol{\theta}}^{(k)}\}_{k\in\{0\}\cup\widehat{\mathcal{C}}}$, threshold parameter $\bar{\lambda}$.
@@
    For $k\in\{0\}\cup\widehat{\mathcal{C}}$, obtain the density estimate $\widehat{f}_{i}^{(k)}$ and conduct the projection (11).
@@
    For $k\in\{0\}\cup\widehat{\mathcal{C}}$, compute $\widehat{\boldsymbol{w}}^{(k)}$ by traditional QR on $\{((\boldsymbol{x}_{i,\widehat{S}^{(k)}}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}$, where the index set $\widehat{S}^{(k)}=\{j:|\widehat{\beta}_{j}+\widetilde{\delta}^{(k)}_{j}|>\bar{\lambda}\}$, then enlarge $\widehat{\boldsymbol{w}}^{(k)}$ (with zero) to $\widetilde{\boldsymbol{w}}^{(k)}\in\mathbb{R}^{p}$.
@@
    Incorporate empirical score functions (12) and solve for
@@ 4.2 Asymptotic Normality for Debiased Estimator
As in Theorem 1, we assume an arbitrarily specified transferable set $\mathcal{C}$. Let $c,C$ be generic constants which can be different at different places.
@@
For each study $k\in\{0\}\cup\mathcal{C}$, (i) $\|\boldsymbol{\theta}_{0}^{(k)}\|_{2}\leq C$; (ii) There exist sparsity levels $\{s_{k}^{*}\}_{k\in\{0\}\cup\mathcal{C}}$ and vectors $\{\boldsymbol{\theta}^{(k)}\}_{k\in\{0\}\cup\mathcal{C}}$ such that $s_{k}^{*}=\max\{s+s_{k},\|\boldsymbol{\theta}^{(k)}\|_{0}\}$, where each $\boldsymbol{\theta}^{(k)}$ satisfies $(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{\theta}_{0}^{(k)}=(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{\theta}^{(k)}+r_{\boldsymbol{\theta}i}^{(k)}$ with $\|\boldsymbol{\theta}_{0}^{(k)}-\boldsymbol{\theta}^{(k)}\|_{1}=Cs_{k}^{*}\sqrt{\log p/n_{k}}$, $\mathbb{E}[(r_{\boldsymbol{\theta}i}^{(k)})^{2}]\leq Cs_{k}^{*}/n_{k}$; (iii) $c\leq\min_{1<j\leq p}\mathbb{E}[(f_{i}^{(k)}x^{(k)}_{i,j}v_{0i}^{(k)})^{2}]^{1/2}\leq\max_{1<j\leq p}\mathbb{E}[(f_{i}^{(k)}x^{(k)}_{i,j}v_{0i}^{(k)})^{3}]^{1/3}\leq C$.
@@
[…]
Conditions 7-9 generalize the assumptions of single-study version in [belloni2019valid] to the transfer setting. Condition 7 set up the moment requirements for approximation error $r_{i}^{(k)}$, where each source’s quantity $r_{i}^{(k)}$ is linked to the whole sample size $n_{\mathcal{C}}$. Condition 8 follows the original settings, providing the necessary requirements for establishing convergence rates of weighted Lasso estimators $\{\widehat{\boldsymbol{\theta}}^{(k)}\}_{k\in\{0\}\cup\mathcal{C}}$. Condition 9 is for the estimation of the conditional density at $(\tau\pm h^{(k)})$-th quantiles, where the approximately sparse QR model is assumed around the quantile index $\tau$.
Remark 6.
Since our inference framework perform debiasing on the transferable source studies, this naturally requires the debiased component across studies to be very similar, which is relatively strict than the similarity requirement for estimation in Theorem 1. As in Condition 7 (iii), each term $\mathbb{E}[f_{i}^{(k)}v_{0i}^{(k)}r_{i}^{(k)}]=\mathbb{E}[(v_{0i}^{(k)})^{2}]\delta_{1}^{(k)}$ needs be sufficient small with order $o(1/\sqrt{n_{\mathcal{C}}})$. Nevertheless, this assumption is reasonable to hold in many practical scenarios. For instance, in multi-source genetic studies, researchers are often interested in estimating the common effect of a specific genetic variant on a phenotype across different sources [michailidou2017association, kunkle2019genetic]. In practice, we further propose a quality control procedure in Section 8.1.4 of supplementary materials. The procedure ensures that if the similarity of the coefficient of interest is not well supported, one should instead debias on the target study only.
@@
Assume Conditions 7-9 and all conditions required by Theorem 1. For each study $k\in\{0\}\cup\mathcal{C}$, let $\lambda_{\tau}^{(k)}=c_{\lambda}\sqrt{n_{k}\log p}$, where $c_{\lambda}$ is a large enough constant. If the parameters satisfy the following growth conditions
$\frac{s\log p}{n_{\mathcal{C}}}+\sum_{k\in\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}\|\boldsymbol{\delta}^{(k)}\|_{1}=o(1/\sqrt{n_{\mathcal{C}}}),$ (13)
$\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\left[\frac{1}{h^{(k)}}\sqrt{\frac{s_{k}^{*}\log p}{n_{k}}}+(h^{(k)})^{2}+\frac{\lambda_{\boldsymbol{\theta}}^{(k)}s_{k}^{*}}{n_{k}}\right]=o(n_{\mathcal{C}}^{-1/4}),$
$\max_{k\in\{0\}\cup\mathcal{C}}\left\{s_{k}^{*}+\frac{n_{k}s_{k}^{*}\log p}{(h^{(k)}\lambda^{(k)}_{\boldsymbol{\theta}})^{2}}+\left(\frac{n_{k}(h^{(k)})^{2}}{\lambda^{(k)}_{\boldsymbol{\theta}}}\right)^{2}\right\}\frac{\log p}{n_{\mathcal{C}}}=o(1/\sqrt{n_{\mathcal{C}}}),$
then for the estimator $\widehat{\alpha}$ obtained from Algorithm 2 with the transferable set $\mathcal{C}$, we have
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\sigma_{\mathcal{C}}^{-1}\sqrt{n_{\mathcal{C}}}(\widehat{\alpha}-\beta_{1}^{*})=\mathbb{U}_{\mathcal{C}}(\tau)+o_{p}(1),\hskip 18.49988pt\mathbb{U}_{\mathcal{C}}(\tau)\leadsto\mathcal{N}(0,1),$
$\text{where}\qquad\mathbb{U}_{\mathcal{C}}(\tau)=\left(\tau(1-\tau)\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[(v_{0i}^{(k)})^{2}]\right)^{-1/2}\cdot\frac{1}{\sqrt{n_{\mathcal{C}}}}\sum_{k\in\{0\}\cup\mathcal{C}}\sum_{i=1}^{n_{k}}\psi_{i}^{(k)}(\alpha,\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)}),$
$\text{and}\qquad\sigma_{\mathcal{C}}^{2}=\left(\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[f_{i}^{(k)}x_{i,1}^{(k)}v_{0i}^{(k)}]\right)^{-2}\cdot\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[\tau(1-\tau)(v_{0i}^{(k)})^{2}].$
@@
Remark 7.
As implied by the minimax lower bound in [li2023transfer], the $\sqrt{n_{\mathcal{C}}}$-rate becomes theoretically possible only when the similarity conditions are linked to the total sample size $n_{\mathcal{C}}$ instead of $n_{0}$ only. As shown in the parameter growth conditions stated in (13), we relate the parameter growth of each individual study to the total sample size $n_{\mathcal{C}}$. Owing to the Neyman orthogonal technique we employ, these requirements are not stringent, where a sufficient condition is to let $h^{(k)}\asymp n_{k}^{-1/6},\lambda_{\boldsymbol{\theta}}^{(k)}\asymp\sqrt{n_{k}\log p},s_{k}^{*}\log p=o(n_{k}^{2/3}/\sqrt{n_{\mathcal{C}}}),\|\boldsymbol{\delta}^{(k)}\|_{1}\log p=o(\sqrt{n_{k}\wedge n_{0}}/\sqrt{n_{\mathcal{C}}})$ for each study $k\in\{0\}\cup\mathcal{C}$.
With the asymptotic normality of $\widehat{\alpha}$ and consistent variance estimators $\widehat{\sigma}_{\mathcal{C}}^{2}$ provided in Theorem 4, we can construct valid confidence intervals and hypothesis test procedures for the objective component $\beta^{*}_{1}$. Now we give some explanations for results in Theorem 4. Firstly, our estimator $\widehat{\alpha}$ enjoys $\sqrt{n_{\mathcal{C}}}$-normality, which is much sharper than the previous results, demonstrating the effectiveness of our debiased transfer methods. Secondly, conditions of Theorem 4 allow for nonidentical covariate distributions across studies, as well as robust to the divergence of covariance matrices between target and source studies, demonstrating our methods’ superiority under covariate shift. Moreover, note that the asymptotic variance $\sigma_{\mathcal{C}}^{2}$ includes $v_{0i}^{(k)}$ that is inversely proportional to the $\tau$-th density $f_{i}^{(k)}$, clearly revealing that residual shift may increase the variance and deteriorate the inference efficiency, and that’s why we choose to only debias on the detected transferable set.